Solve for x using the addition principle.
x + 16 = -8



I have to admit that i am at a total loss. When i decided to take online classes i didnt know algebra would be one of them. Please help me. I am too old for this lol.

Answer by stanbon(75887)

x + 16 = -8
Add -16 to both sides of the equation to get:
x+16+(-16) = -8+(-16)
x + 0 = -24
x = -24
==============
Cheers,
Stan H.